# Affordance-based Robot Manipulation with Flow Matching

### v3

@@ 3.2.2 Flow Matching for Visuomotor Policy Learning
The visual embeddings $\bm{o}$ are obtained through ResNet (He et al., 2016). We also evaluate the Transformer structure in our experiments. Various types of inputs include state-based inputs, RGB images, and visual affordances.
Closed-loop action trajectory prediction: We execute the action trajectory prediction obtained by our flow matching model for a fixed duration before replanning. At each step, the policy takes the observation data $\bm{o}$ as input and predicts $Tp$ steps of actions, of which $Ta$ steps of actions are executed on the robot without re-planning. $Tp$ is the action prediction horizon and $Ta$ is the action execution horizon. The policy predicts flow vectors $\mathbf{v}_{t}$ conditioned on visual observation data $\bm{o}$ with Feature-wise Linear Modulation (FiLM) (Perez et al., 2018) as well as the interpolated waypoints $\bm{x}_{t}$. The flow model $f_{\bm{\theta}}$ is represented with U-Net (Ronneberger et al., 2015). The whole training process of the flow matching policy is illustrated in Algorithm 1. In the subsequent experiments, we will further study multiple model variants, including transformer-based structure, trajectory horizon, etc.
In our case of robot manipulation, $\bm{x}_{1}$ in Equation (3) represents the demonstration robot action trajectories. $\bm{x}_{0}$ is the random generated waypoints following a multivariate normal distribution $\bm{x}_{0}\sim\mathcal{N}(0,I)$. $\bm{x}$ here could denote 6D robot end-effector trajectories or robot joint actions.
Inference: For the inference procedure, random waypoints are sampled from the source distribution and then flowed into the target trajectory by estimating the flow from $t=0$ to $t=1$ over steps:
@@
3.2.3 Intriguing Properties of Flow Matching Policy
In this section, we provide some insights and intuitions about flow matching policy and its comparisons against diffusion policy for clarification. Algorithm 1 and Algorithm 2 respectively shows the pseudocode of training flow matching and diffusion policy with Denoising Diffusion Probabilistic Models (DDPM). We could see several differences between these two methods:
  1. 1.
@@ 3.2.3 Intriguing Properties of Flow Matching Policy
  3. 3.
    Flow matching also includes a particularly interesting family of probability paths: the vector field that corresponds to the principle of Optimal Transport (OT) with linear interpolant. Paths with linear interpolation are simpler than diffusion paths, forming straight line trajectories whereas diffusion paths result in curved paths. These properties seem to empirically translate to more stable training, faster generation, and better performance.
Algorithm 1 Robot Flow Matching Policy Training
observation $\bm{o}$, target robot actions $\bm{x}_{1}$, source random waypoints $p_{0}$
flow $\mathbf{v}_{\bm{\theta}}$
while not converged do
     $\bm{x}_{0}\sim p_{0}$, sample random robot waypoints
     $t\sim\mathcal{U}[0,1]$, sample time steps
     $\bm{x}_{t}=t\bm{x}_{1}+(1-t)\bm{x}_{0}$, linear interpolation
     $\mathbf{v}_{t}(\bm{x}|\bm{o})=v_{\bm{\theta}}(\bm{x}_{t},t|\bm{o})$, flow estimation
     $\nabla_{\bm{\theta}}\left\|v_{\bm{\theta}}(\bm{x}_{t},t|\bm{o})-\dot{\bm{x}}_{t}\right\|$, gradient step
end while
Stopping criteria: $1,000$ training epochs
@@
observation $\bm{o}$, target robot actions $\bm{x}_{1}$, source Gaussian noises $p_{0}$
noise $\mathbf{\epsilon}_{\bm{\theta}}$
while not converged do
     $\bm{x}_{0}\sim p_{0}$, sample Gaussian noises
     $t\sim\mathcal{U}[0,1]$, sample time steps
     $\bm{x}_{t}=\mathcal{N}(\mathbf{x}_{t};\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I})$, forward process
     $\mathbf{\epsilon}_{t}(\bm{x}|\bm{o})=\epsilon_{\bm{\theta}}(\bm{x}_{t},t|\bm{o})$, noise estimation
     $\nabla_{\bm{\theta}}\left\|\epsilon_{\bm{\theta}}(\bm{x}_{t},t|\bm{o})-\bm{\epsilon}_{t}\right\|$, gradient step
end while
Stopping criteria: $1,000$ training epochs
3.3 Activities of Daily Living Dataset
We construct a real-world dataset with $10$ tasks across Activities of Daily Living to test the proposed method. Each task includes $1,000$ sets of RGB images, demonstrated robot trajectories, and labeled ground truth of affordances. Thus $10,000$ demonstrations have been collected in total. The data has been manually collected by moving robot end-effectors with kinesthetic teaching. We label the heatmap affordance of our data with 2D Gaussian blobs centered on the pixel of demonstrated action during data collection. The novelty of our dataset could be summarized as: (i) It contains the same scenarios with multiple objects, multi-task affordances, and the demonstrated robot trajectories. (ii) All tasks are related to Activities of Daily Living that involve (simulated) human data.
Objects are randomly placed within the range of the table (1.5 meters $\times$ 1 meter). The human (manikin) is placed randomly around the table in the camera view. We have around 30 different objects. The camera view is fixed as we have not used 3D representation, which will be addressed in our future work.
Our tasks include prompt primitives: ‘sweep the trash’, ‘pass the water to the human’, ‘hang the towel’, ‘put on the hat’, ‘cover the food’, ‘wipe the nose’, ‘wipe the forearm’, ‘feed the human’, ‘comb the hair’, and ‘brush the teeth’. Each task corresponds to one prompt primitive. We could also add a pretrained LLM layer (e. g., GPT) in the front for zero-shot text classification, allowing for linking other language instructions to one of the ten prompt primitives. The dataset will be made publicly available soon.
4 Experiments
We systematically evaluate prompt tuning method and flow matching policy against baseline studies. We also investigate how design choices would affect their performance. We randomly split the Activities of Daily Living dataset with $80$%-$20$% percentage of training and testing. The results reported here are obtained after $1,000$ epochs of training with 3-fold cross validation to minimize the impact of any noises.
4.1 Affordance Evaluation with Prompt Tuning
4.1.1 Baseline Studies
In this experiment, we benchmark our proposed shallow and deep prompt tuning structures against several commonly used finetuning and instruction-aware vision encoding protocols.
@@ 4.1.1 Baseline Studies
  • Adapter-based methods: insert MLP layers with residual connections between pretrained frozen transformer layers of vision and language, as customary in the literature (Gao et al., 2024).
  • Decoder-based methods: These methods treat the pretrained backbone as a feature extractor with fixed weights during tuning, and only the decoder is tuned, as customary in the literature (He et al., 2016).
  • side-network methods: train a text transformer network on the side and append pretrained vision features and sidetuned text features before being fed into the decoder, as customary in the literature (Ganz et al., 2024).
  • Cross-attention methods: Similarly to the above side-network methods, the difference here is using cross-attention fusing instead of simple prepending. An example of cross-attention fusing vision and language could be found in the literature (Jiang et al., 2022).
For a fair comparison, all the baselines here use self-supervised pretrained MAE weights on ImageNet-21k dataset for the vision transformer model. In ablation studies, we will investigate different pretrained weights. The training parameters for the prompt tuning network include an image size of 224 × 224, AdamW with a learning rate of 1.5e-4 including Warmup with step-decay, and batch size of 256. The text encoding layers output 76 tokens, as the CLIP method (Radford et al., 2021).
4.1.2 Main Results
@@
Table 1: Results of prompt tuning baseline and ablation studies. We report the number of learnable parameters, the heatmap estimation error (the fourth column) and the heatmap center error (the fifth column). Our method outshines other baselines except for the full finetuning.
Table 1 presents the results of prompt tuning on our testing dataset for affordance learning, comparing against baselines. We use two metrics to evaluate our results: (i) L2 error of affordance heatmap estimation (the fourth column), and (ii) L2 distance between the predicted and ground truth of heatmap centers (the fifth column). For the inference, we fit Gaussian Mixture Models to determine the heatmap centers. The heatmap error is averaged on each map, and the center error is averaged on per center point.
Four observations could be made from this result:
  1. 1.
    The deep structure outperforms the other baselines except for the full finetuning.
  2. 2.
    Full finetuning slightly outperforms deep prompt tuning in terms of heatmap estimation error and heatmap center error. However, the distinction of heatmap center errors ($1.78$ pixels) remains subtle, given the full image size of $224\times 224$. This outcome is favorable as it indicates that most heatmap errors are caused by the tails of the Gaussian distribution, instead of the center area where the robot actions actually applied on.
  3. 3.
    As expected, the shallow structure performs suboptimally to the deep structure. In the following ablation studies, we will investigate how the prompt depth affects the performance.
  4. 4.
    Generalizability: We also observe that the trained model could be generalized to new objects. For example, the training dataset only includes a manikin. We found out that it generates well on our testing data with real humans. Note that the proposed tuning method is parameter-efficient, it is envisaged that the method could be readily transferred to different tasks with a small amount of task-specific data.
@@
What do prompts learn? We show a t-SNE (Van der Maaten and Hinton, 2008) visualization of the embeddings after the last vision transformer layer and before the decoder in Fig. 5. We can see that the points of the same color (e. g., tasks with the same language prompts) are embedded together, implying that the representations recover the underlying manifold structure of discriminative task information.
@@
4.1.3 Ablations
We further ablate model design choices.
Pretrained Weights. We evaluate using MAE self-supervised pretrained weights and supervised pretrained weights trained on ImageNet-21k dataset for the vision transformer model. The results in Table 1 show self-supervised pretrained weights perform better.
We are aware of other variants of vision transformers or weights, for example, CLIP vision encoder. As our goal is to integrate textual representations into any vision encoder while keeping it frozen, preserving its visual understanding capabilities, we have chosen the most basic ViT-B-16 transformer backbone and commonly used pretrained weights. We have also tried GPT directly, but no constant performance was achieved without carefully designed prompts or finetuning.
Decoder Input. We apply the decoder on the global output and image-corresponding output after the vision transformer respectively and report results in Table 1.
Dataset Size. We use various amounts of data to train the model. Fig. 6-left shows that prompt tuning has better adaptability than full finetuning when downstream data is scarce.
Prompt Location. We have seen different conclusions from prior works about whether the vision-language fusion should be integrated at early or late transformer layers. Thus we conduct experiments to insert prompts at various layers. From Fig. 6-right, we can see that inserting prompts to early layers (for example, layer 1-3 from bottom to top) achieves higher loss than inserting to late layers (for example, layer 1-3 from top to bottom). Thus in our case, prompts have greater significance at the late transformer layers. These results are also supported by the nature of vision transformer hierarchy: lower layers mainly capture low-level fundamental visual details, while higher layers focus on high-level concepts that might be vital for downstream tasks.
@@ 4.1.3 Ablations
4.2 Flow Matching Policy Evaluation
4.2.1 Baseline Studies
We compare our flow matching policy against two other robot behavior cloning methods: (i) Diffusion Policy (Chi et al., 2023) with Denoising Diffusion Probabilistic Models (DDPM), and (ii) Transformer-based behavior cloning with Mean Square Error Loss, as customary in RVT (Goyal et al., 2023), RT-X (Padalkar et al., 2023).
Note that we are aware of other well-performed robot behavior cloning methods, including energy-based IBC (Florence et al., 2022), GAIL (Ho and Ermon, 2016), etc. Since extensive studies have been conducted and showed lower performance of these methods compared to diffusion policy, we choose two representative baselines for evaluation.
We consider CNN-based and Transformer-based structures for flow matching. The flow matching and baselines are trained in a supervised manner using RGB images with visual affordances as input, and the output is a long-horizon trajectory in both 2D pixel space and 3D Cartesian space. All 10 tasks in our dataset are trained in a single policy. Table 4 shows the hyperparameters we have used in flow matching and diffusion policy. Table 5 shows the tasks Summary.
@@ 4.2.1 Baseline Studies
Baselines Diffusion Policy (2-step inference) 38.37 25.08 24.81
Diffusion Policy (16-step inference) 0.884 2.096 159.72
Transformer-based BC 2.797 6.109 7.59
Ours Flow Matching (Transformer) 2.151 4.842 18.05
Flow Matching (CNN, 1-step inference) 0.898 2.151 8.53
Ablations Flow Matching (CNN, 2-step inference) 0.888 2.104 24.29
Flow Matching (CNN, 4-step inference) 0.846 2.094 44.99
Flow Matching (CNN, 8-step inference) 0.842 2.092 70.10
Flow Matching (CNN, 16-step inference) $0.841$ $2.091$ 100.98
Flow Matching (task-specific policy) 1.107 3.41 9.19
@@
Table 2: Results of flow matching policy baseline and ablation studies. We report the error of 2D and 3D trajectory estimation and the average inference time. Our method achieves the best trajectory estimation accuracy, compared to other baselines. We also investigate the effect of various design choices on flow matching performance, including network structure and inference steps.
4.2.2 Main Results
Table 2 presents the results of flow matching policy on our testing dataset for robot trajectory learning, comparing against baselines. We use two metrics for evaluation: (i) the error of 2D and 3D trajectory estimation, and (ii) the average inference time across various steps, performed using PyTorch with RTX 4090 GPU acceleration. The trajectory error is averaged on each point of the trajectory.
@@ 4.2.2 Main Results
Figure 7: Training and testing loss of flow matching and diffusion policy throughout the training process. Flow matching exhibits greater stability on both training and evaluation than diffusion policy.
Three observations could be made from this result:
  1. 1.

### v4

@@ III-B2 Flow Matching for Visuomotor Policy Learning
Algorithm 1 Robot Flow Matching Policy Training
1:observation $\bm{o}$, target robot actions $\bm{x}_{1}$, source random waypoints $p_{0}$
2:flow $\mathbf{v}_{\bm{\theta}}$
3:while not converged do
4:     $\bm{x}_{0}\sim p_{0}$, sample random robot waypoints
5:     $t\sim\mathcal{U}[0,1]$, sample time steps
6:     $\bm{x}_{t}=t\bm{x}_{1}+(1-t)\bm{x}_{0}$, linear interpolation
7:     $\mathbf{v}_{t}(\bm{x}|\bm{o})=v_{\bm{\theta}}(\bm{x}_{t},t|\bm{o})$, flow estimation
8:     $\nabla_{\bm{\theta}}\left\|v_{\bm{\theta}}(\bm{x}_{t},t|\bm{o})-\dot{\bm{x}}_{t}\right\|$, gradient step
9:end while
10:Stopping criteria: training epochs reached
Closed-loop action trajectory prediction: We execute the action trajectory prediction obtained by our flow matching model for a fixed duration before replanning. At each step, the policy takes the observation data $\bm{o}$ as input and predicts $Tp$ steps of actions, of which $Ta$ steps of actions are executed on the robot without re-planning. $Tp$ is the action prediction horizon and $Ta$ is the action execution horizon. The whole training process of flow matching is illustrated in Algorithm 1.
Inference: For the inference procedure, random waypoints are sampled from the source distribution and then flowed into the target trajectory by estimating the flow from $t=0$ to $t=1$ over steps. We could adopt multiple steps $1/\Delta t$ for inference:
@@
III-B3 Implementation Details
For the network structures of flow matching, we first adopt ResNet [16] for visual embeddings $\bm{o}$. The flow model $f_{\bm{\theta}}$ is represented with U-Net [40]. The flow model predicts vectors $\mathbf{v}_{t}$ conditioned on visual observation embeddings $\bm{o}$ with Feature-wise Linear Modulation (FiLM) [36] as well as the interpolated waypoints $\bm{x}_{t}$. In the subsequent experiments, we will further study multiple model variants, including transformer-based structure, trajectory representation, etc.
In our case of robot manipulation, $\bm{x}_{1}$ in Equation (3) represents the demonstration robot action trajectories. $\bm{x}_{0}$ is the random generated waypoints following a multivariate normal distribution $\bm{x}_{0}\sim\mathcal{N}(0,I)$. $\bm{x}$ here could denote 6D robot end-effector trajectories, robot joint actions, gripper actions, etc. The visual embeddings $\bm{o}$ include various types of inputs, such as state-based inputs, RGB-D images, and visual affordances.
III-B4 Comparisons against Diffusion Policy
In this section, we provide some insights and intuitions about flow matching and its comparisons against diffusion policy for clarification. Algorithm 1 and Algorithm 2 respectively shows the pseudocode of training flow matching and diffusion policy with DDPM. We can see several differences between these two methods:
@@ III-B4 Comparisons against Diffusion Policy
  • Flow matching includes a particularly interesting family of probability paths: the vector field that corresponds to the principle of Optimal Transport with linear interpolant. Flow matching paths with linear interpolation are simpler than diffusion paths, forming straight-line trajectories whereas diffusion paths result in curved paths. These properties seem to empirically translate to more stable training, faster generation, and better performance.
@@
1:observation $\bm{o}$, target robot actions $\bm{x}_{1}$, source Gaussian noises $p_{0}$
2:noise $\mathbf{\epsilon}_{\bm{\theta}}$
3:while not converged do
4:     $\bm{x}_{0}\sim p_{0}$, sample Gaussian noises
5:     $t\sim\mathcal{U}[0,1]$, sample time steps
6:     $\bm{x}_{t}=\mathcal{N}(\mathbf{x}_{t};\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I})$, forward process
7:     $\mathbf{\epsilon}_{t}(\bm{x}|\bm{o})=\epsilon_{\bm{\theta}}(\bm{x}_{t},t|\bm{o})$, noise estimation
8:     $\nabla_{\bm{\theta}}\left\|\epsilon_{\bm{\theta}}(\bm{x}_{t},t|\bm{o})-\bm{\epsilon}_{t}\right\|$, gradient step
9:end while
10:Stopping criteria: training epochs reached
III-C Activities of Daily Living Dataset
We construct a real-world dataset with $10$ tasks across Activities of Daily Living. Each task includes $1,000$ sets of RGB-D images, demonstrated robot action trajectories, and labeled ground truth of affordance maps. Thus $10,000$ demonstrations have been collected in total. The data has been manually collected by moving robot end-effectors with kinesthetic teaching. The novelty of our dataset includes: (i) It contains the same scenarios with multiple objects, multi-task affordances, and the demonstrated robot trajectories. (ii) All tasks are related to Activities of Daily Living that involve (simulated) human data.
We label the affordance heatmaps with 2D Gaussian blobs centered on the object pixels of the demonstrated action. The affordance maps model the locations of all relevant object areas that physically interact with robots given each task. For example, ’feeding task’ requires affordance heatmaps centered on the fork handle, food, and human mouth.
Objects are randomly placed within the range of the table (1.5 meters $\times$ 1 meter). The human (manikin) is placed randomly around the table in the camera view. We have around 30 different objects. Our tasks include prompt primitives: ‘sweep the trash’, ‘pass the water to the human’, ‘hang the towel’, ‘put on the hat’, ‘cover the food’, ‘wipe the nose’, ‘wipe the forearm’, ‘feed the human’, ‘comb the hair’, and ‘brush the teeth’. We can also add a pretrained LLM layer (e. g., GPT) in the very front for zero-shot text classification, allowing for linking other language instructions to one of the ten prompt primitives. The camera is positioned with some variations but generally oriented towards the table and objects. Please refer to the supplementary materials for more videos of each task.
IV Experiments
We systematically evaluate the proposed prompt tuning and flow matching methods against baseline studies. We also ablate how design choices would affect their performance.
IV-A Affordance Evaluation with Prompt Tuning
IV-A1 Baseline Studies
We benchmark our proposed prompt tuning structures against several commonly used finetuning and instruction-aware vision encoding protocols:
@@ IV-A1 Baseline Studies
  • Adapter-based methods: insert MLP layers with residual connections between pretrained frozen transformer layers of vision and language, as customary in the literature [10].
  • Decoder-based methods: These methods treat the pretrained backbone as a feature extractor with fixed weights during tuning, and only the decoder is tuned, as customary in the literature [16].
  • Side-network methods: train a text transformer network on the side and append pretrained vision features and sidetuned text features before being fed into the decoder, as customary in the literature [9].
  • Cross-attention methods: Similarly to the above side-network methods, the difference here is using cross-attention fusing instead of simple prepending. An example of cross-attention fusing vision and language could be found in the literature [25].
For a fair comparison, all the baselines here use self-supervised pretrained MAE weights on ImageNet-21k dataset for the vision transformer model. We randomly split our Real-world Activities of Daily Living (ADLs) dataset with $80$%-$20$% percentage of training and testing. The results reported here are obtained after $1,000$ epochs of training.
IV-A2 Main Results
@@
TABLE I: Results of prompt tuning baseline and ablation studies. We report the number of learnable parameters, the heatmap estimation error (the fourth column) and the heatmap center error (the fifth column). Our method outshines other baselines except for the full finetuning.
Table I presents the results of prompt tuning on our ADLs testing dataset for affordance learning, comparing against baselines. We use two metrics to evaluate our results: (i) L2 error of affordance heatmap estimation, and (ii) L2 distance between the predicted and ground truth of heatmap centers. We fit Gaussian Mixture Models on predicted heatmaps to determine the inferred heatmap centers. The heatmap error is averaged on each map, and the center error is averaged on per center point. Three observations could be made:
  • Prompt tuning against full finetuning The deep structure of prompt tuning outperforms other baselines except for full finetuning. Full finetuning slightly outperforms deep prompt tuning in terms of heatmap estimation error and heatmap center error. However, the distinction of heatmap center errors ($1.78$ pixels) remains subtle, given the full image size of $224\times 224$. This outcome is favorable as it indicates that most heatmap errors are caused by the tails of the Gaussian distribution, instead of the center area where the robot actions actually applied on. We will further ablate the impact of dataset size on these two methods.
  • Generalizability: We also observe that the trained model could be generalized to new objects. For example, the training dataset only includes a manikin. We found out that it generates well on our testing data with real humans. Affordances on objects with similar shapes (e. g., forks and spoons) could also be transferred. Note that as the proposed tuning method is parameter-efficient, it is envisaged that the method could be readily transferred to different tasks with a small amount of task-specific data.
  • What do prompts learn? We show a t-SNE [46] visualization of the embeddings after the last vision transformer layer (before the decoder) in Fig. 4. We can see that the points of the same color (e. g., tasks with the same language prompts) are embedded together, implying that the representations recover the underlying manifold structure of discriminative task information.
@@
IV-A3 Ablations
We further ablate model design choices:
Pretrained Weights: We evaluate using MAE self-supervised pretrained weights and supervised pretrained weights trained on ImageNet-21k dataset for the vision transformer model. The results in Table I show self-supervised pretrained weights perform better. We are aware of other more complicated variants of vision transformers, for example, CLIP vision encoder and its pretrained weights. As our goal is to integrate textual representations into any vision encoder while keeping it frozen, we have chosen the most basic ViT-B-16 transformer backbone and commonly used pretrained weights and achieved competitive results. We have also tried GPT directly to infer affordance, but no constant performance was achieved without carefully designed prompts or finetuning.
Decoder Input: We apply the decoder on the global output and image-corresponding output after the vision transformer respectively and report results in Table I.
Dataset Size: We use various amounts of data for training. Fig. 5-left shows that prompt tuning has better adaptability than full finetuning when downstream data is scarce.
Prompt Location: We have seen different conclusions from prior works about whether the vision-language fusion should be integrated at early or late transformer layers. We conduct experiments to insert prompts at various layers. From Fig. 5-right, we can see that inserting prompts to early layers (for example, layer 1-3 from bottom to top) achieves higher loss than inserting to late layers (for example, layer 1-3 from top to bottom). Thus in our case, prompts have greater significance at the late transformer layers. These results are also supported by the nature of vision transformer hierarchy: lower layers mainly capture low-level fundamental visual details, while higher layers focus on high-level concepts that might be vital for downstream tasks.
@@ IV-A3 Ablations
IV-B Flow Matching Policy Evaluation
IV-B1 Baseline Studies
We compare our flow matching policy against: (i) Diffusion Policy [3] with DDPM and DDIM, and (ii) Transformer-based behavior cloning, as customary in RVT [13], RT-X [35].
Some previous research [34, 11] has mathematically proven that DDIM trajectories are equivalent to flow matching trajectories with a different scaling of time. This scale difference may be beneficial for flow matching. In this experiment, we aim to practically evaluate such speculation. Table V shows the hyperparameters used in flow matching and diffusion policy.
Note that we are aware of other competitive robot behavior cloning methods, including energy-based IBC [7], GAIL [18], etc. Since extensive studies have been conducted and showed better performance of the diffusion policy against these methods, we choose the representative transformer and diffusion baselines for evaluation.
We first train and test flow matching and baselines on our collected Real-world Activities of Daily Living (ADLs) dataset in a supervised manner with Mean Square Error Loss. In the following sections, we will also evaluate on other benchmarks (e. g., Push-T, Franka Kitchen, and Robomimic).
  • Real-world Activities of Daily Living (ADLs) dataset: For 2D data, the training uses a RGB image with visual affordances as input, and the output is a trajectory in 2D pixel space. For the counterparts in 3D, the training takes the concatenation of RGB-D images with visual affordances as input, and outputs a trajectory in 3D Cartesian space. We randomly split the dataset with $80$%-$20$% percentage of training and testing. The results reported are obtained after $1,000$ epochs of training. For each baseline, all 10 tasks in our dataset are trained in a single policy in a supervised manner. To ensure fair comparisons, the corresponding hyperparameters across the flow matching policy and baseline methods are selected to remain consistent. Table VI shows the task summary.
@@ IV-B1 Baseline Studies
Ours Flow Matching (Transformer) 2.151 4.677 18.05
Flow Matching (CNN, 16-step) $0.840$ $1.009$ 98.981
Ablations Flow Matching (CNN, 1-step) 0.888 1.031 13.228
Flow Matching (CNN, 4-step) 0.846 1.014 41.494
Flow Matching (CNN, 8-step) 0.842 1.013 78.222
Flow Matching (without affordance) 1.107 1.908 13.450
Baselines DDPM (1-step) 2.851 2.479 13.791
DDPM (4-step) 0.890 2.411 45.736
DDPM (8-step) 0.884 2.403 80.200
DDPM (16-step) 0.882 2.398 99.197
DDIM (1-step) 4.328 10.78 13.757
DDIM (4-step) 0.876 2.299 42.672
DDIM (8-step) 0.875 2.272 80.940
DDIM (16-step) 0.874 2.266 98.680
Transformer-based BC 2.797 4.911 7.59
@@
TABLE II: Results of flow matching policy against baselines and ablations. We report the average error of 2D and 3D trajectory estimation and the inference time. Our flow matching method achieves the best trajectory estimation accuracy. We also investigate the effect of various design choices on flow matching performance, including network structures, training, and inference steps.
@@
Figure 6: Training and testing loss throughout the training process. Flow matching exhibits greater stability on both training and evaluation than diffusion policy.
IV-B2 Main Results
Table II presents the results of flow matching policy on our ADLs testing dataset for robot trajectory learning, comparing against baselines. We use two metrics for evaluation: (i) the error of 2D and 3D trajectory estimation, and (ii) the inference time, performed with RTX 4090 GPU. The trajectory error is averaged on each point of the trajectory.
Four observations could be made from this result: